7) How many moles are in 125 grams of oxygen difluoride, OF2?

(O = 16.00 g/mol, F = 19.00 g/mol)

Group of answer choices

54.0

2.31

0.43

67.5

8) Which of the following contains the same number of moles of atoms as 6.00 grams of carbon?

Group of answer choices:

12.00 g of Si

12.15 g of Mg

6.00 g of Si

8.00 g of He

9) Which sample would contain the largest number of atoms?

(C = 12.01 g/mol, Si = 28.08 g/mol, S = 32.1 g/mol, Au = 196.97 g/mol)

Group of answer choices:  

19.6 g of Au

8.0 g of S

7.0 g of Si

4.0 g of C

10) How many grams would it take to make 0.25 moles of carbon disulfide, CS2?

(C= 12.01 g/mol, S = 32.1 g/mol)

Group of answer choices

14.0

19.0

76.2

11.0
Answer #1

7)

mass of OF2 = 125 g

moles of OF2 = 125 / 54 = 2.31

Answer : 2.31

8) Which of the following contains the same number of moles of atoms as 6.00 grams of carbon?

mass of C = 6.00 g

moles of C = 6.00 / 12 = 0.500 mol

moles of Mg = 12.15 / 24.3 = 0.500 mol

answer : 12.15 g of Mg

9)

moles of Au = 19.6 / 196.97 = 0.0995

moles of S = 8 / 32 = 0.25

moles of Si = 7 / 28 = 0.249

moles of C = 4 / 12 = 0.33

carbon has largest number of atoms

Answer : 4.0 g of C

10)

(C= 12.01 g/mol, S = 32.1 g/mol)

moles of CS2 = 0.25

mass of CS2 = 0.25 x 76.21 = 19.0 g

Answer : 19.0